OpenStudy (anonymous):

Brett has determined a function f(x) that shows the exponential growth of the number of shoes Larae owns each year. Explain how the f^−1(x) can be found and what f^−1(132) means. @shamil98 @nincompoop

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@FutureMathProfessor

OpenStudy (anonymous):

@FutureMathProfessor plss help meee

OpenStudy (shamil98):

f(x) . \[f^{-1}(x)\] is the inverse function of f(x), it can be found by replacing the x and y values like if f(x) = x+2 y = x+ 2 inverse is x = y + 2 x - 2 = y x-2 = f^-1(x)

OpenStudy (shamil98):

replacing the values and then solving for y*

OpenStudy (anonymous):

then what would be next?

OpenStudy (shamil98):

f^{-1)(132) you input 132 into the x value(s) of the inverse function and evaluate it. then you would plot the points on a graph where x is 132 and y is the result of 132 input in the function